The theoretical probability of undesirable side effects resulting from taking Grebex is 2 in 13. If 260 people take Grebex to lower their blood pressure, how many will encounter undesirable side effects?

Found 2 solutions by stanbon, rfer:
Answer by stanbon(75887)

The theoretical probability of undesirable side effects resulting from taking Grebex is 2 in 13. If 260 people take Grebex to lower their blood pressure, how many will encounter undesirable side effects?
----------------------------------------
2 to 13 is the same as 2x to 13x
-----
Equation:
2x+3x = 260
5x = 260
x = 52
----
2x = 104 will encounter undesirable side effects.
----------------
Cheers,
Stan H.

Answer by rfer(16322)

2/13=x/260
13x=520
x=40